OpenStudy (anonymous):

If f(x) = x^2 + 2x + 3, what is the average rate of change of f(x) over the interval [-4, 6]?

OpenStudy (anonymous):

(A).51 (B). 40 (C). 31 (D). 20 (E). 4

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{ find f(6) and f(-4) then } \hspace{.33em}\\~\\ \normalsize \text{ avg rate of change }=\dfrac{f(6)-f(-4)}{6-(-4)} \hspace{.33em}\\~\\\dfrac{}{} }\end{align}\)

OpenStudy (anonymous):



OpenStudy (anonymous):